41 F equals how many degrees Celsius?

How many cubic feet are in 3 cubic yards?

Answer by elima(1433)

41 F equals how many degrees Celsius?
The formula would be:
C=5/9(F-32)
C=5/9(41-32)
C=5/9(9)
c=45/9
C=5
How many cubic feet are in 3 cubic yards?
There are 3 feet to 1 yard so you would multiply 3*3=9
So there are 9 cubic feet to 3 cubic yards.
=)